# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v2

@@ 2.1.4. Our Contributions
Our study incorporates elements of human-AI co-creation, crowd creativity, and collective dynamics. We note that co-creation studies often confound the effect of exposure with the effect of disclosure: If one is creating with an AI system, it is impossible to separate the content of an AI system from the knowledge that the content is from an AI system. Our factorial design lets us estimate the marginal effect of exposure and disclosure separately. Co-creation studies typically employ a small number of specialized participants actively engaged with a system. From the perspective of validating a system, this is reasonable. But we are interested in the effects of (1) passive exposure on (2) a general public. For this reason, we adopt a large-scale experimental design—similar to crowd-sourced creativity studies—that lets us estimate effects on the general public rather than specialized users. A key benefit of our large sample size is that we can precisely estimate how participant attitudes affect human-AI outcomes. This is important because, as Gero et al. (2022, pg. 1016) write: “[P]articipant attitudes are a major unknown factor when studying human-AI collaboration.” Drawing on the “many-worlds” paradigm, our experiment design also lets us understand the effect of AI over time since responses feed forward, allowing us to observe differences in rates of change between conditions.
2.2. Factor 1: LLM Exposure
2.2.1. Effects on Creativity
Intuitively, the effect of exposure to ChatGPT ideas will depend on how creative ChatGPT answers are relative to human ideas. In preliminary testing, we found that the answers to the AUT generated by our prompt were scored as more creative than the ideas generated by humans (see Appendix E) via the Organisciak et al. (2022) classifier. LLM generations may be increasing in creativity: while GPT-3 (an earlier model than ChatGPT-3.5) scored lower in AUT creativity than humans on the AUT (Stevenson et al., 2022), GPT-4 (a more recent model than ChatGPT-3.5) scored among the top percentile of humans on a similar verbal creativity task (Guzik et al., 2023), as measured by human judges.
2.2.2. Effects on Diversity
2.3. Factor 2: LLM Disclosure
2.3.1. Effects on Creativity
2.3.2. Effects on Diversity
It is not clear how knowing content is from AI will affect the diversity of ideas participants produce. But prior work suggests heterogeneity along two lines: the difficulty of the prompt [^4] As discussed later, we measure the ‘difficulty’ of a prompt by the inverse rank of the average creativity in the control condition. If participants tended to submit lower creativity ideas in the control condition for item X, we said item X was difficult., and the attitude of the participant. Prior work suggests that disclosing ideas as AI-generated would decrease diversity due to automation bias, the tendency to over-rely on AI systems (Schemmer et al., 2022; Mosier et al., 1996; Goddard et al., 2014). Increased reliance on AI ideas (when labeled as such) could lead to lower idea diversity and higher AI adoption. Conversely, some evidence suggests people display algorithmic aversion to creative products such as haikus (Hitsuwari et al., 2022) or art (Kirk et al., 2009). This aversion would yield the opposite prediction. Roemmele (2021) found that seeing AI examples only affected the participant’s writing on a key measure for difficult prompts—suggesting creative task difficulty might moderate the effect of disclosure on AI adoption. Task confidence decreases reliance on automated systems and trust in a system increases reliance on automated systems (Goddard et al., 2014). Although this literature is not usually applied to creativity, we might then suspect that people self-reporting low creativity (i.e., low task confidence) and those who think AI is more creative than humans (i.e., high system trust) are most likely to increase adoption of AI ideas when the source is disclosed.
3. Pre-Experiment
Before describing the experiment, we discuss how we chose the five specific AUT (Alternate Uses Test) items and constructed our ChatGPT prompt.
3.1. Stimuli Construction
3.1.1. Choosing AUT Items
We had to choose a selection of items that people would brainstorm creative uses for. We chose five items for which the creativity classifier that we used had the highest accuracy. Previously, Organisciak et al. (2022) fine-tuned GPT-3 Davinci to predict the creativity of AUT items. This dataset contains 20,121 responses from 2,025 participants, across 21 distinct AUT items and nine distinct studies (Organisciak et al., 2022). [^5] We obtained this dataset by direct correspondence with Dr. Organisciak on February 23, 2023; the code that Dr. Organisciak used to generate this dataset is available at https://github.com/massivetexts/llm_aut_study/blob/main/notebooks/Process_AUT_GT.ipynb Each response was graded for creativity by humans and normalized to a scale of 1-5. Then Organisciak et al. (2022) fine-tuned GPT-3 Davinci on this dataset to predict creativity scores. Here, fine-tuning involves providing {Input (an AUT response), Output (human rating)} pairs to a pre-trained LLM. Then the LLM adjusts its parameters to produce a similar output given an input, proxying human judgments. Overall, the fine-tuned GPT-3 classifier had a correlation [^6] We obtained scores for this classifier by downloading the zip file from (https://github.com/massivetexts/llm_aut_study/blob/main/results/evaluation.zip), then navigating to gt_main2/gpt-ft-davinci-1.csv of $r=0.81$ with human judgment. Accuracy varied by item (Appendix Table LABEL:aut_desc_stats). For our experiment, we picked the five items for which the classifier had the highest accuracy: tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is ‘in-domain’ for the classifier since we ask participants to do the same task for the same items the classifier was trained on.
3.1.2. Generating GPT Ideas
We generated AI ideas with ChatGPT-3.5 and a zero-shot prompt based on prior work. These decisions followed two principles: ecological validity and precedent.
Model & Prompting Strategy
Our model and prompting strategy were driven by a desire to approximate how ordinary users would use large language models to generate ideas. First, we used ChatGPT-3.5, the latest ChatGPT model freely available at the time of the study. Because ChatGPT has a popular and accessible UI, we assume users would be more likely to use ChatGPT rather than a model available only through an API or on a limited basis. Second, we used zero-shot prompting rather than few-shot prompting. Because zero-shot prompting requires no labeled data, this would be a more natural use case for most users.
Prompt Construction
Our specific zero-shot prompt was informed by prior work on LLMs and creativity. Stevenson et al. (2022) administered the AUT to GPT-3 through a zero-shot prompt. However, this prompt generated much wordier responses $(M=25.4,SD=8.5)$ than the human responses in the Organisciak Dataset $(M=4.6,SD=5)$. Such a discrepancy would alert participants to what was AI vs. human generated, which would nullify the disclosure factor (whether the source of an idea is disclosed). Hence, we appended a request (Figure 2) to use roughly the same number of words (5) as the average human response. The modified prompt resulted in responses with an average word length $(M=4.4,SD=1.3)$ much closer to human responses than the un-modified Stevenson prompt, $p<0.001$ by two-tailed permutation tests. We use the ideas from this modified Stevenson prompt for our experiments (Appendix B for more details).
Figure 2. To generate AUT ideas, we used the zero-shot prompt from Stevenson et al. (2022) with an additional instruction at the end to match the mean length of human responses from prior work.
@@ 4.1. Summary
We recruited participants from a mixture of social media and newsletters. Once participants clicked on the link to the experiment, they were taken to a landing page. In addition to a consent button, that landing page asked several questions. Participants were asked (1) to rate their creativity relative to other humans and to AI, (2) to rate their attitude towards AI (Nadeem, 2022, 2023), (3) age, (4) country, and (5) gender. After consenting, participants engaged in 5 trials. For each trial, a participant generated a creative use for an item under a specific experimental condition, after seeing example ideas. These example ideas constituted our experimental manipulation. Ideas fed forward to future trials such that if a participant was in the {[Control], tire} condition the example ideas the participant saw were the most recent ideas from prior participants in the {[Control], tire} condition. See Table 1 for experimental conditions. The experiment took place in the summer of 2023.
Table 1. Experiment conditions and associated factors. Participants complete the Alternate Uses Task in each condition after being exposed to prior responses generated in each condition. Conditions vary by LLM exposure (none, high or low) and LLM disclosure (source is labeled or not).
Condition Number of AI Number of Human Source Disclosed
Examples Examples (Y/N)
Control 0 6 N
High Exposure; Disclosed 4 2 Y
High Exposure; Not Disclosed 4 2 N
Low Exposure; Disclosed 2 4 Y
Low Exposure; Not Disclosed 2 4 N
4.2. Ethics
The experiment was approved by our university’s institutional review board.
4.3. Recruitment
@@
We recruited volunteer participants through three sources: (1) Facebook ads, (2) Reddit, and (3) the weekly newsletter of Creative Mornings [^7] https://creativemornings.com/ [^8] The first author contacted Creative Mornings, who agreed to include the experiment in the newsletter., which is ‘the world’s largest face-to-face creative community’. Creative Mornings is an organization geared towards creative professionals that organizes (e.g.) talks and meetups. We ensured all participants were above 18. While we did not offer monetary compensation, we offered to give participants information about themselves, such as their creativity relative to both humans and AI and their ability to spot creative ideas. Providing information to participants about themselves is often effective for recruiting volunteer participants since it makes the task intrinsically rewarding (Reinecke and Gajos, 2015). Appendix F describes the information we provided to participants.
@@
We recruited volunteer participants instead of crowdsourced workers for several reasons. First, we wanted participants to be intrinsically motivated since (1) many theories suggest intrinsic motivation helps creativity (Mumford and Hemlin, 2017) and (2) we did not want low-quality engagement to confound results (especially since ideas propagate forward). Second, we were interested in an international sample. Because we did not pay participants, we did not need to collect any personally identifiable information. Each user was assigned a random identifier. The experiment being anonymous created a lower barrier to recruiting international participants since GDPR was not operative. Finally, we recruited participants in a targeted manner. In particular, we wanted to generalize this experiment to two key groups: individuals who have a demonstrated interest in technology and those who have a demonstrated interest in creativity. These groups are most relevant to the phenomena in question. To this end, we reached technology-oriented users by posting the experiment in the following subreddits: r/InternetIsBeautiful, r/chatgpt, r/singularity, and r/artifical. We reached creativity-oriented users by posting the experiment in r/writing, r/poetry, and the Creative Mornings newsletter. We also used several ‘neutral’ sources to test the experiment: r/samplesize and Facebook ads. If a participant completed the experiment, then the participant was given a shareable link to their results so they could spread the study.
@@ 4.4. Experiment Procedure
Once participants clicked on our link, they were taken to a landing page that included a consent form, task description, and pre-treatment questions.
4.4.1. Study Description
The description read as follows:
What you will do:
We’ll show you 5 common items, and you’ll come up with creative uses for each item. To spark your imagination, you’ll see ideas from prior participants and even from AI (i.e., ChatGPT). You’ll be asked to rank these ideas in order of creativity. The ideas you write may be anonymously shown to future participants to spark their imagination. The study takes 3-6 minutes to complete. The goal is to learn about how humans and AI brainstorm.
What you will learn:
  • How creative you are compared to other humans
  • How creative you are compared to AI
  • How well you can rank creative ideas
@@ 4.4.1. Study Description
See Appendix F for more details on how each of these three pieces of information was calculated.
4.4.2. Pre-Treatment Questions
Participants were asked several pre-treatment questions:
  1. (1)
    (required) A slider ranging from 0 to 100 that says ‘I am more creative than Xof AI‘
  2. (2)
    (required) A slider ranging from 0 to 100 that says ‘I am more creative than Xof Humans‘
  3. (3)
    (required) ‘Artificial intelligence computer programs are designed to learn tasks that humans typically do. Would you say the increased use of artificial intelligence computer programs in daily life makes you feel…[‘More concerned than excited’, ‘More excited than concerned’, ‘Equally excited and concerned’]
  4. (4)
    (optional) What country are you from?
  5. (5)
    (optional) What is your age?
  6. (6)
    (optional) What is your gender?
The third question was from Pew (Nadeem, 2023, 2022). Our gender question was based on guidance from Spiel et al. (2019). We chose the Pew question instead of a longer battery of questions about AI to minimize the response burden. See Appendix C for more details about these questions.
@@ 4.4.4. Task Instructions
For each trial, participants were asked to first rank a list of example ideas in order of creativity and then submit their own idea:
Task
For this task, you will submit a creative use for a [ITEM]. But before submitting your idea, here are some ideas for inspiration. Rank them by creativity.
Rank Previous Ideas
  • Rank these ideas in order of creativity, with the most creative use on top. Drag ideas to rank them.
  • We’ll show you how your rankings compare to rankings from a highly accurate model.
@@
Submit Your Idea
Your turn! What is a creative use for a [ITEM]? The goal is to come up with a creative idea, which is an idea that strikes people as clever, unusual, interesting, uncommon, humorous, innovative, or different. List a creative use for a [ITEM].
See Appendix M for screenshots. We asked participants to rank ideas to ensure that they would engage with the example ideas. [^9] We did not use these rankings as a DV since—because a participant ranks the examples they are shown, and all examples are from the same condition—these ranks could not speak to between-condition differences, which is the focus of the paper. Depending on the condition, (1) either there were or were not AI ideas in this example set (exposure); (2) AI ideas were or were not labeled (disclosure). We use the same prompt for humans (the text under Submit Your Idea) as with ChatGPT (Stevenson et al., 2022), but with a slight modification to request a single idea. This prompt contains language consistent with best practices in divergent thinking assessment (Beaty et al., 2021). After submitting an idea, participants received feedback on their idea’s uniqueness and how accurately they ranked the example ideas (Appendix M).
@@ 4.4.5. Response Chains
Note that because human ideas are propagated at the {[condition], item} level, the human ideas in the [Control] condition are ‘clean’ from AI contamination. They were brainstormed after seeing sets of human-only ideas, also from the [Control] condition.)
[…]
@@
5. Recruited Participants
@@
Unique Countries 48.00
Total Responses 3414.00
Unique Participants 844.00

### v3

@@ 2.1.3. Our Contributions
Our study integrates human-AI co-creation, crowd creativity, and collective dynamics. Traditional co-creation studies typically confound exposure and disclosure effects: when creating with AI, one cannot separate the AI’s content from the knowledge that this content is AI-generated. Our design allows us to estimate these effects independently. Co-creation studies typically employ a small number of specialized participants actively engaged with a system. We are interested in the effects of (1) passive exposure on (2) a more general public. For this reason, we adopt a large-scale experimental design—similar to crowd-sourced creativity studies—that lets us estimate effects on the general public rather than specialized users. A key benefit of our large sample size is that we can precisely estimate how participant attitudes affect human-AI outcomes. This is important because attitudes affect human-AI collaboration outcomes (Gero et al., 2022). Drawing on the “multiple-worlds” paradigm (Salganik and Watts, 2009), our experiment design also lets us understand the effect of AI over time since responses feed forward, allowing us to observe differences in rates of change between conditions. This method can be used for other experiments to measure the cumulative effect of AI.
@@ 3.1. Summary
We recruited participants through social media platforms and newsletters. Once participants clicked on the link to the experiment, they were taken to a landing page. In addition to a consent button, that landing page asked several questions related to their self-perceived creativity, attitudes towards AI, and demographics. Following consent, participants completed 5 trials. Each trial required generating a creative use for an object under a specific experimental condition after viewing example ideas—which were our experimental manipulation. Ideas fed forward to future trials such that if a participant was in the {[Control], tire} condition the example ideas the participant saw were the most recent ideas from prior participants in the {[Control], tire} condition. See Figure 1 for experimental conditions. The experiment was approved by our University’s IRB.
3.2. Stimuli Creation
Picking AUT Items.
Previously, Organisciak et al. (2023) fine-tuned GPT-3 to predict the creativity of AUT items. Since we use this classifier for our creativity DV, we picked the five items for which the classifier had the highest classifier-human judgment correlations: tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is highly ‘in-domain’ for the classifier: Participants do the same task, using the same items, from the classifier training and evaluation. See Appendix A.1 for more details on the classifier dataset (“Organisciak Dataset”), training, and performance.
Model & Prompting Strategy.
We generated AI ideas using ChatGPT-3.5 with a zero-shot prompt adapted from (Stevenson et al., 2022). We chose ChatGPT-3.5 for ecological validity as it was the freely available model powering ChatGPT at the time of our study. Zero-shot prompting required no labeled data, making it more practical for typical users. We modified the prompt from Stevenson et al. (2022) to match word counts from the Organisciak Dataset (Organisciak et al., 2023), preventing easy identification of AI-generated content. Using parameters from (Stevenson et al., 2022) and instructions to match human word counts, our modified prompt produced responses similar to human responses in length. These generated ideas served as stimuli for our main experiment. For prompt details and statistical tests, see Appendix A.2.
3.3. Recruitment Strategy
@@
We recruited volunteer participants from three sources: (1) Facebook ads, (2) Reddit, and (3) Creative Mornings’ weekly newsletter. Creative Mornings [^4] https://www.creativemornings.com/, ”the world’s largest face-to-face creative community,” is an organization for creative professionals that hosts talks and meetups. Creative Mornings agreed to distribute our experiment. We verified all participants were at least 18 years old. While we did not offer monetary compensation, we offered to give participants information about themselves, such as their creativity relative to both humans and AI, and their ability to spot creative ideas. Providing information to participants about themselves is often effective for recruiting volunteers since it makes the task intrinsically rewarding (Reinecke and Gajos, 2015). See Table 3 for sources and SI 6 for sample characteristics.
@@
We recruited volunteer participants instead of crowdsourced workers for several reasons. First, we wanted participants to be intrinsically motivated since (1) many theories suggest intrinsic motivation helps creativity (Mumford and Hemlin, 2017) and (2) we did not want low-quality engagement to confound results (especially since ideas propagate forward). Second, we recruited participants in a targeted manner. In particular, we wanted to target (1) individuals who have a demonstrated interest in technology and (2) those who have a demonstrated interest in creativity. These groups are most relevant to the phenomenon in question. We reached technology-oriented users by posting the experiment in the following subreddits: r/chatgpt, r/InternetIsBeautiful, r/singularity, and r/artificial. We reached creativity-oriented users by posting the experiment in r/writing, r/poetry, and the Creative Mornings newsletter. We also used several ‘neutral’ sources to test the experiment: r/samplesize and Facebook ads. If a participant completed the experiment, then the participant was given a shareable link to their results so they could spread the study.
@@ 3.4. Experiment Procedure
Once participants clicked on our link, they went to a landing page that included a consent form, task description, and pre-treatment questions. See Figure 3 for the task description. See SI 5 for more details on how participant feedback was calculated.
What you will do: We’ll show you 5 common items, and you’ll come up with creative uses for each item. To spark your imagination, you’ll see ideas from prior participants and even from AI (i.e., ChatGPT). You’ll be asked to rank these ideas in order of creativity. The ideas you write may be anonymously shown to future participants to spark their imagination. The study takes 3-6 minutes to complete. The goal is to learn about how humans and AI brainstorm.
What you will learn: How creative you are compared to other humans How creative you are compared to AI How well you can rank creative ideas
@@
Figure 3. Study description provided to participants.
3.4.1. Pre-Treatment Questions.
Participants were asked several pre-treatment questions (Figure 4), including self-perceived creativity, perceived creativity of AI, and attitude towards AI. The third question was from Pew (Nadeem, 2023, 2022). We chose the Pew question instead of a longer battery of questions about AI to minimize the response burden. See SI 2 for more details about these questions.
(1) (required) A slider ranging from 0 to 100 that says ‘I am more creative than Xof AI‘ (2) (required) A slider ranging from 0 to 100 that says ‘I am more creative than Xof Humans‘ (3) (required) ‘Artificial intelligence computer programs are designed to learn tasks that humans typically do. Would you say the increased use of artificial intelligence computer programs in daily life makes you feel…[‘More concerned than excited’, ‘More excited than concerned’, ‘Equally excited and concerned’] (4) (optional) What country are you from? (5) (optional) What is your age? (6) (optional) What is your gender?
Figure 4. Pre-Treatment Questions
@@ 3.4.3. Task Instructions
For each trial (Figure 5), we asked participants to first rank a list of example ideas by creativity and then submit their own idea. See Appendix D for screenshots. We included the ranking task to ensure engagement with the example ideas. Across conditions, we manipulated two factors: (1) exposure—whether AI-generated ideas were included in the example set, and (2) disclosure—whether AI ideas were labeled. We used the same prompt for human participants as Stevenson et al. (2022) used with ChatGPT, with a minor modification requesting a single idea (for reasons discussed in 1.2.1). This prompt incorporates language aligned with divergent thinking assessment best practices (Beaty et al., 2021). After idea submission, participants received feedback on their idea’s uniqueness and their ranking accuracy (screenshots in Appendix D).
Task For this task, you will submit a creative use for a [ITEM]. But before submitting your idea, here are some ideas for inspiration. Rank them by creativity.
Rank Previous Ideas Rank these ideas in order of creativity, with the most creative use on top. Drag ideas to rank them. We’ll show you how your rankings compare to rankings from a highly accurate model.
@@
Submit Your Idea Your turn! What is a creative use for a [ITEM]? The goal is to come up with a creative idea, which is an idea that strikes people as clever, unusual, interesting, uncommon, humorous, innovative, or different. List a creative use for a [ITEM].
Figure 5. Task Instructions
@@ 3.4.4. Response Chains
Note that because human ideas are propagated at the {[condition], item} level, the human ideas in the [Control] condition are ‘clean’ from AI contamination. They were brainstormed after seeing sets of human-only ideas, also from the [Control] condition.
We ran seven response chains for each of the 25 (5 items $\times$ 5 conditions) combinations, corresponding to 175 response chains in all and 3500 targeted responses (175 response chains $\times$ 20 trials per chain).
@@
4. Outcome Measures
Table 1. Three outcome measures (idea diversity, creativity, AI adoption) across three levels of analysis (local, global, evolution). All ideas embedded using SBERT.
Local Global Evolution
Creativity How creative is the submitted response? Measured by the prediction from the Organisciak et al. classifier. Not applicable Does the creativity of submitted ideas change over time? Measured by the slope of trial number (i.e., iteration in response chain) on creativity scores from Organisciak et al.
Idea Diversity How different is a participant’s response from example responses? Measured by median pairwise semantic distance between ideas a participant sees and their response (SI 8.3). How diverse were all participants’ ideas in a condition? Measured by median pairwise distance between all submitted ideas in a condition. Do ideas become more different from each other as the experiment progresses? We measure the median pairwise distance of submitted ideas at each trial number, then calculate the slope of trial number on idea diversity.
AI Adoption How similar is a participant’s response to AI example responses? Measured by the maximum semantic distance between a participant’s response and AI examples. Not applicable Not applicable
We have three outcome measures (idea diversity, creativity, and AI adoption) and three levels of analysis (local, evolution, and global). See Table 1. The local level measures outcomes at the level of an individual trial (e.g., how a submitted response relates to example responses). The evolution level measures the rate of change of outcome variables with respect to the trial number in the response chain (i.e., experiment iteration). The global level compares all submitted responses in a condition to each other. For all pairwise comparisons, we use a Holm-Bonferroni adjustment for multiple comparisons. For idea diversity and AI adoption, we scale the dependent variable (cosine distance or cosine similarity, respectively) by 100 for easier interpretation. See Appendix B for more methodological details.
5. Results
5.1. Recruited Participants
We received over 3,000 responses from 48 countries. See SI 6 for sample characteristics. Out of a total of five trials, participants finished four trials on average (Table 2), suggesting the experiment was engaging. Most participants came from the Creative Mornings newsletter or r/InternetIsBeautiful (Table 3 for source counts and categorization). Of the non-missing responses, the sample was 50% women, 43% men, 4% non-binary, 3% not disclosed, 1% self-described. The mean age was 34.92 (SD = 10.86). Regarding AI, the sample was 48% neutral, 28% excited, 24% concerned. Participants said they were more creative than 57.86% (SD = 26.66) of AI and 58.67% (SD = 23.65) of humans. See SI 6 for kernel density plots.
@@ 5.1. Recruited Participants
Unique Countries 48
Total Responses 3414
Unique Participants 844